OpenStudy (anonymous):

find the solution of m(-13 - m)=0

OpenStudy (ilovecake):

What do ya think?

OpenStudy (ilovecake):

Any ideas?

OpenStudy (anonymous):

idk lol thats why i came here

OpenStudy (anonymous):

0 and 13???

OpenStudy (anonymous):

@ilovecake

OpenStudy (ilovecake):

yes

OpenStudy (anonymous):

uhm WRONG

OpenStudy (ilovecake):

That is correct.

OpenStudy (anonymous):

no i just selected that and it wasnt

OpenStudy (ilovecake):

what? I took my test that had the same question and I got 100%

OpenStudy (mathstudent55):

\(m(-13 - m) = 0\) \(m = 0\) or \(-13 - m = 0\) \(m = 0\) is correct from the left equation above. Now solve the right equation above. What do you get for m? You need to solve \(-13 - m = 0\)

OpenStudy (mathstudent55):

The second solution is not 13.